# A Comprehensive Study on Large Language Models for Mutation Testing

### v3

@@ Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we report the results of a comprehensive empirical study over six different LLMs, including both state-of-the-art open- and closed-source models, on 851 real bugs drawn from two different Java real-world bug benchmarks. Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with higher fault detection. That is, 79.1% (for LLMs) vs. 41.6% (for rule-based); an increase of percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points respectively. These findings are immediately actionable for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
@@ 1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In those most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Foster et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021) reported on results from deployment of mutation testing at their companies. In particular, results recently reported by Meta (Foster et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
@@
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, there is to date, no large-scale comprehensive empirical study. Without such a study, we have no basis on which to deploy an industrial scale with confidence. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
The generation of mutants, using any technique is challenging because of the following three important issues:
Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
@@
This paper reports on how well Large-Language Models (LLMs) handle these challenges, compared to more traditional rule-based approaches. We conducted a comprehensive comparison between six LLMs (three closed-source commercial and three open-source ones) and four existing mutation testing techniques with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our experiments required the generation of 611,100 mutations.
@@
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency). Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
@@
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This leads to a higher fault detection rate than traditional approaches. For example, the mutants generated by GPT-4o-Mini lead to a fault detection rate of 90.8%, while the mutants generated by LEAM, PIT, and Major are 63.9%, 40.1%, and 66.9%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations. Moreover, the distributions of AST edit distance of mutations compared to the original code also present that LLMs generate more balanced and expressive mutations. For instance, almost 70% of the mutations of Major have the AST edit distance of one, while for GPT-4o, only 50% of its mutations have the same distance.
@@
We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
@@
We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
@@
We build a high-quality dataset of Java mutations comprehensively labeled with detailed quality and behavior metrics.
@@ 2.1. Overview and Research Questions
RQ3 (Efficiency) What are the time costs of different mutation generation approaches? This research question investigates the time costs of different mutation generation approaches. Time costs are associated with the efficiency and practicality of mutation generation approaches.
@@ 2.2. Studied Mutation Generation Approaches
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches is for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. Table 1 summarizes the mutation generation approaches covered by this study.
@@
| LLM-based | GPT-3.5 | GPT-3.5-Turbo, whose based model is GPT, trained on data before 2021/09, released in 2021/03. |
| GPT-4o | The based model is GPT, trained on data before 2023/10, released in 2024/05. | |
| GPT-4o-M | GPT-4o-Mini, whose based model is GPT, trained on data before 2023/10, released in 2024/07. | |
| SC-16b | StarChat--16b, whose based model is StarCoder, released in 2023/06. | |
| CL-13b | CodeLlama-Instruct-13b, whose based model is Llama, released in 2023/08. | |
| DS-236b | DeepSeek-Coder-V2-236b, whose based model is DeepSeek, released in 2024/07. |
@@ 2.2.1. Traditional Approaches
LEAM: LEAM (Tian et al., 2022) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
@@ 2.2.2. LLM-based Approaches
Our study involve multiple kinds of LLMs, including both open-source and closed-source models.
@@
LLMs: Our study also comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), and DS-236b (DeepSeek-Coder-V2-236b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.
@@ 2.3. Datasets
To this end, we employ the Defects4J 2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 2. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total. However, from Table 1 and Table 2, we observe the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder222https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.

### v4

@@ Abstract.
Large Language Models (LLMs) have recently been used to generate mutants in both research work and in industrial practice. However, there has been no comprehensive empirical study of their performance for this increasingly important LLM-based Software Engineering application. To address this, we conduct a comprehensive empirical study evaluating BugFarm and LLMorpheus (the two state-of-the-art LLM-based approaches), alongside seven LLMs using our newly designed prompt, including both leading open- and closed-source models, on 851 real bugs from two Java real-world bug benchmarks. Our results reveal that, compared to existing rule-based approaches, LLMs generate more diverse mutants, that are behaviorally closer to real bugs and, most importantly, with higher fault detection. That is, 87.98% (for LLMs) vs. 41.64% (for rule-based); an increase of percentage points. Nevertheless, our results also reveal that these impressive results for improved effectiveness come at a cost: the LLM-generated mutants have worse non-compilability, duplication, and equivalent mutant rates by , , and percentage points, respectively. These findings are immediately actionable for both research and practice. They allow practitioners to have greater confidence in deploying LLM-based mutation, while researchers now have a baseline for the state-of-the-art, with which they can research techniques to further improve effectiveness and reduce cost.
@@ 1. Introduction
For many years, mutation testing has been the topic of scientific research (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). Nevertheless, until the last five years, it had made little penetration into the industrial sector. In the most recent five years, it has become increasingly widely deployed in industry. For example, large tech sector companies such as Meta (Harman et al., 2025; Beller et al., 2021) and Google (Petrović et al., 2021b; Petrović and Ivanković, 2018) reported on results from the deployment of mutation testing at their companies. In particular, results recently reported by Meta (Harman et al., 2025) highlighted the pivotal role played by Large Language Models (LLMs) in the successful deployment of mutation testing at scale in industry.
@@
However, despite increasing interest in LLM-based mutation testing in both research and practitioner communities, to date, no large-scale comprehensive empirical study. Without such a study, we have no basis on which to deploy an industrial scale with confidence. Furthermore, the research community currently has no reliable baseline results for the state-of-the-art effectiveness and efficiency of LLM-based mutation testing compared to more traditional rule-based mutation generation techniques (Jia and Harman, 2010; Papadakis et al., 2019).
The generation of mutants, using any technique, is challenging because of the following three important issues:
Generating effective mutations: The ultimate goal of mutation testing is to lead to strong tests that uncover real faults, making it critical that generated mutants closely couple with real faults (Chekam et al., 2020; Wen et al., 2019; Petrović et al., 2021a). While existing methods have shown promise, there is still significant room for improvement in enhancing their fault-revealing potential (Degiovanni and Papadakis, 2022).
@@
This paper examines how well Large-Language Models (LLMs) address these challenges, in comparison to more traditional rule-based approaches. We conducted a comprehensive empirical study which studies BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025) (i.e., the state-of-the-art LLM-based mutation generation approaches), alongside seven LLMs adopting our newly designed prompt, covering both leading open- and closed-source models, and four existing mutation testing techniques (i.e., Major (Just et al., 2011), PIT (Coles et al., 2016), LEAM (Tian et al., 2022), and BERT (Degiovanni and Papadakis, 2022)) with 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014b) and 246 bugs from ConDefects (Wu et al., 2024)). Overall, our primary experiments required the generation of 666,148 mutations.
@@
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of behavior similarity with real bugs (i.e., real bug detection rate, average Ochiai coefficient score, the number of high-Ochiai score bugs, coupling rate, and coupling categorization) as well as mutant validity (i.e., compilability rate, duplicate mutation rate, equivalent mutation rate, and efficiency). Additionally, we study how different prompt engineering strategies, turbulence on prompt, and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), DeepSeek-V3 (Liu et al., 2024), StarChat--16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. We also explore the potential of using LLMs for generating high-order mutations. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
@@
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This results in a higher fault detection rate compared to traditional approaches. For instance, the mutants generated by DeepSeek-V3-761b and GPT-4o-Mini achieve fault detection rates of 91.1 % and 90.8%, respectively, whereas those generated by LEAM, PIT, and Major yield substantially lower rates of 63.9%, 40.1%, and 66.9%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, duplicate mutation rate, and equivalent mutation rate. In particular, Major excels with a 97.6% compilability rate, 0% duplicate mutations, and 2.0% equivalent mutants, while GPT-4o exhibits a compilability rate of 76.4%, a duplicate mutation rate of 7.8%, and an equivalent mutation rate of 3.1%, respectively.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that the DeepSeek models exhibit the greatest diversity, newly introducing 49 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 21. The existing LLM-based approach, BugFarm, introduces 31 node types, higher than the traditional approaches as well. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations. Moreover, the distributions of AST edit distance of mutations compared to the original code also show that LLMs generate more balanced and expressive mutations. For instance, almost 70% of the mutations of Major have an AST edit distance of one, whereas for GPT-4o, only 50% of its mutations have the same distance.
@@
We also categorise the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall into 9 compilation error types, with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. Moreover, we further analyze the code context of the non-compilable mutations generated by LLM-based approaches, and find that Method Reference and Method Invocation are the most frequent code features. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, thereby enhancing the effectiveness of mutation testing using AI models.
We analyze the alive (i.e., survive) mutations of each approach, which are important to strengthen the test case. For one alive non-equivalent mutation, it can be (1) not covered, (2) covered but no state infection, or (3) state infection without detection. We analyze the distribution of alive mutations across different approaches and find that LLMs based on our prompt are more likely to generate uncovered mutations, which consequently remain alive, providing more valuable guidance for enhancing unit tests.
@@
We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our study not only covers other existing LLM-based mutation generation approaches, but also introduces a new effective prompt template. Our findings indicate that LLMs excel in generating diverse mutations that closely mimic real bugs.
@@
We construct a high-quality dataset of Java mutations, comprehensively annotated with detailed quality and behavior metrics, which can also serve as a valuable resource for fine-tuning in future research.
@@ 2.1. Overview and Research Questions
RQ3 (Efficiency) What are the general costs of different mutation generation approaches? This research question investigates the general costs of different mutation generation approaches. We consider time and token costs in this RQ, which are associated with the efficiency and practicality of mutation generation approaches.
@@
RQ5 (Alive Mutations) What are the distributions of alive mutations across different approaches? This RQ aims to investigate the underlying reasons why mutations remain alive, by categorizing them into three cases following the existing study (Just et al., 2014a): (1) uncovered by tests, (2) covered but without infecting the program states, and (3) injecting the states but remaining undetected by tests. By examining how these cases are distributed across different approaches, we can better understand the effectiveness of generated mutations and their potential value for guiding test suite improvement.
@@ 2.2. Studied Mutation Generation Approaches
Our study is conducted in Java because: 1) it is one of the most widely used programming languages, shown as the TIOBE ranking111https://www.tiobe.com/tiobe-index/; 2) the majority of existing mutation generation approaches are for Java. We categorize these approaches into two groups: traditional approaches (i.e., those developed before the emergence of LLMs) and LLM-based approaches (i.e., approaches that leverage LLMs and are designed for interaction in human-like conversations), as shown below. The details of the settings for each generation approach are illustrated in Section 2.5. Table 1 summarizes the mutation generation approaches covered by this study.
@@
| LLM-based | BF (GPT-4o) | The existing LLM-based approach BugFarm with GPT-4o. |
| LLMph (GPT-4o) | The existing LLM-based mutation generation approach LLMorpheus with GPT-4o. | |
| LLMut (GPT-3.5) | Our prompt with GPT-3.5-Turbo, trained on data before 2021/09, released in 2021/03. | |
| LLMut (GPT-4o) | Our prompt with GPT-4o, trained on data before 2023/10, released in 2024/05. | |
| LLMut (GPT-4o-M) | Our prompt with GPT-4o-Mini, trained on data before 2023/10, released in 2024/07. | |
| LLMut (SC-16b) | Our prompt with StarChat--16b, whose base model is StarCoder, released in 2023/06. | |
| LLMut (CL-13b) | Our prompt with CodeLlama-Instruct-13b, whose base model is Llama, released in 2023/08. | |
| LLMut (DS-236b) | Our prompt with DeepSeek-Coder-V2-236b, whose base model is DeepSeek, released in 2024/07. | |
| LLMut (DS-671b) | Our approach with DeepSeek-V3-761b, whose base model is DeepSeek, released in 2024/12. |
@@ 2.2.1. Traditional Approaches
LEAM: LEAM (Tian et al., 2022, [n. d.]) is the state-of-the-art custom-trained-model-based approach, which employs sequence-to-sequence models to generate mutations. Its model is trained on a large corpus of real-world bugs and their corresponding patches, allowing it to learn patterns for generating more effective mutations.
@@ 2.2.2. LLM-based Approaches
Our study incorporates two existing LLM-based mutation generation approaches (i.e., BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025)), and introduces a new prompt design for mutation generation, which will be presented later.
Existing LLM-based Approaches: We adopt two the state-of-the-art LLM-based mutation generation approaches, BugFarm (Ibrahimzada et al., 2023) and LLMorpheus (Tip et al., 2025). BugFarm trains a small-scale model for selecting lines for mutation, and then utilizes LLMs to mutate the target line. LLMorpheus performs mutation given the whole method as context. Note that LLMorpheus was originally designed for JavaScript. Upon examining its prompt, we found it to be language-agnostic. Therefore, we reimplemented the code-extraction component for Java (such as parsing ASTs and extracting the code context for mutation) while retaining the original prompt. For both approaches, we employ GPT-4o (Achiam et al., 2023), the current most widely used model, as their backend LLM.
LLMs with Our Prompt: Different from the existing LLM-based approaches, we propose a new prompt engineering for mutation generation, shown as Figure 1. For ease of presentation, in the following part of this paper, we call our prompt LLMut.
To craft effective prompts, we follow the best practices, which suggest that prompts should comprise four aspects: Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. The placeholder MUT_NUMBER specifies the number of mutations required, while the placeholder CODE_ELEMENT specifies the target code snippet to be mutated. In the Context, we clearly state that a mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs, enabling LLMs to learn how to mutate code from real bugs. We further explore the different prompt settings in Section 4.1 and the impact of prompt turbulence in Section 4.2.
To avoid the few-shot examples leaking information, beyond the evaluation datasets (i.e., as shown later, they are Defects4J and ConDefects), we employ another benchmark, QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 2. As LLM capabilities continue to advance, the optimal number of few-shot examples may vary. We therefore investigate the impact of different numbers of few-shot examples in Section 4.1.3.
| Correct Version | Buggy Version |
| n = (n & (n - 1)); | n = (n ˆ (n - 1)); |
| while (!queue.isEmpty()) | while (true) |
| return depth==0; | return true; |
ArrayList r = new ArrayList();
r.add(first).addll(subset); to_add(r); |
to_add.addAll(subset); |
| c = bin_op.apply(b,a); | c = bin_op.apply(a,b); |
Compared with the prompts used in BugFarm222https://github.com/Intelligent-CAT-Lab/BugFarm and LLMorpheus333https://github.com/githubnext/llmorpheus, our prompt design offers several advantages:
it incorporates real-world bugs as few-shot examples, which have been shown to significantly enhance LLM performance in many code-related tasks;
it determines the number of generated mutants based on the size of the target code, rather than on a fixed number;
it specifies the mutation location using the target statement itself rather than line numbers, enabling more precise localization (Fan et al., 2023a); and
it employs JSON output formatting, which is widely used in LLM fine-tuning and downstream tasks.
@@
Based on LLMut, our study comprehensively covers the most widely studied LLMs. Specifically, we select a representative set of LLMs with code capabilities, both from commercial closed-source and open-source models. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from the GPT family, including GPT-3.5 (GPT-3.5-Turbo) (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-M (GPT-4o-Mini) (Achiam et al., 2023)), and open-source LLMs, including SC-16b (StarChat--16b), CL-13b (CodeLlama-Instruct-13b), DS-236b (DeepSeek-Coder-V2-236b), and DS-671b (DeepSeek-V3-671b). The details of these models are shown as Table 1. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 671 billion. Moreover, the selected LLMs represent a balance between general-purpose models (e.g., GPT-4o and DeepSeek-V3-671b), which are trained for a wide range of natural language and reasoning tasks, and code-specialized models (e.g., CodeLlama-Instruct-13b and DeepSeek-V2-Coder-236b), which are further fine-tuned on large-scale code corpora. This selection allows us to evaluate mutation generation across both broadly capable models and those explicitly optimized for programming tasks, ensuring the conclusions are representative of the current landscape of LLMs.
@@ 2.3. Datasets
To this end, we employ the Defects4J 2.0 (Just et al., 2014b) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 3. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014c; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains historical bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from Defects4J 2.0, which contain 605 bugs in total. However, from Table 1 and Table 3, we observe that the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder444https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs. Among them, only the most recent model, DeepSeek-V3-731b, cannot guarantee data leakage-free, since it is released after the ConDefects dataset.